OpenStudy (anonymous):

calculate the mass in grams in 10 moles of H2O

OpenStudy (superhelp101):

Well, H2O is 18 grams per mole.

OpenStudy (superhelp101):

i guess if you multiply it by 10 like 10X18. idk

OpenStudy (superhelp101):

i am here lol

OpenStudy (anonymous):

Hye

OpenStudy (somy):

mole= mass/Mr mass= mole* Mr

OpenStudy (somy):

@superhelp101 you are right btw Mr is molecular mass

OpenStudy (unklerhaukus):

\[M_\text H=1.01[\text g /\text{mol}]\\ M_\text O=16.00[\text g /\text{mol}]\\~\\ M_{\text{H}_2\text O}=2M_\text H+M_\text O\\ \qquad\quad=(2\times1.01+16.00)[\text g /\text{mol}]\\ \qquad\quad=18.02[\text g /\text{mol}]\] \[m=n\times M\\ \quad=10[\text{mol}]\times18.02[\text g /\text{mol}]\\ \quad=180.2[\text g]\]